$19.30 13.90 83.80 18.90 105.00 S 2,170 Quality control 35 1,500 1,500 3,000 and round your final answers to the nearest whole dollar amount
Unit Cos costing system (Round your Select the appropriate cost driver for each cost pool and calculate the activity rates if Kelier wanted to implement an ABC system, (Round your answers to 2 decimal places)
whole dollar Unit Cost
media%2Fefb%2Fefb8c35f-c97c-438d-95d4-5b
Answer #1

Question 1.

Under traditional costing system, the overheads are assigned to unit based on labour hours, machine hours, units of production, etc. In this example, it is mentioned that the overheads will be assigned on the basis of machine hours.

Therefore, we first need to calculate the overheads per machine hour.

Overhead per machine hour = Total overheads / Total machine hours

Overhead per machine hour = 11,560/3,000 = $3.85333 per hour

Total overheads assigned to Sandy Beach Model = 1500*3.8533 = $5,780

Total overheads assigned to Rocky River Model = 1500*3.8533 = $5,780

Total overheads = 5,780 + 5,780 = 11,560

Question 2.

Overhead cost per unit for Sandy Beach Model = 5,780/1,220 = $4.74

Overhead cost per unit for Rocky River Model = 5,780/920 = $6.28

Total cost per unit = Direct materials + Direct labour + Overheads

Total cost per unit (Sandy Beach) = 19.30 + 13.90 + 4.74 = $37.94

Total cost per unit (Rocky River) = 26.80 + 18.90 + 6.28 = $51.98

Question 3.

Gross margin per unit = Selling price per unit - Production cost per unit

Gross margin per unit (Sandy Beach) = 83.80 - 37.94 = $45.86

Gross margin per unit (Rocky River) = 105.00 - 51.98 = $53.02

Question 4.

Cost driver is the factor that drives the cost. In this example, Set-up cost is the cost incurred due to number of set ups. The quality control cost is incurred due to number of inspections. The maintenance is due to use of the machine.

Activity rate is calculated as cost pool divided by the quantity of cost driver.

Cost centers for each cost pool and activity rates are as below:

Cost driver of set up cost is number of set ups. Activity rate for set-up costs = 2,170/35 = $62 per set up

Cost driver of quality control is number of inspections. Activity rate for quality control = 6,300/525 = $12 per inspection

Cost driver of maintenance is number of machine hours. Activity rate for maintenance = 3,090/3,000 = $1.03 per hour

Question 5.

Overheads cost assigned to each product as follows:

Overhead cost = Cost driver quantity of product*Activity rate

Overhead cost for Sandy Beach

Set-up cost = 20*62 = $1,240

Quality control = 120*12 = $1,440

Maintenance = 1,500*$1.03 = $1,545

Total overhead cost = $4,225

Overhead cost for Rocky River

Set-up cost = 15*62 = $930

Quality control = 405*12 = $4,860

Maintenance = 1,500*$1.03 = $1,545

Total overhead cost = $7,335

Question 6.

Overheads cost per unit = Total overheads / Total units

Overheads cost per unit (Sandy Beach) = 4,225/1,220 = $3.46

Overheads cost per unit (Rocky River) = 7,335/920 = $7.97

Total cost per unit = Direct materials + Direct labour + Overheads

Total cost per unit (Sandy Beach) = 19.30 + 13.90 + 3.46 = $36.66

Total cost per unit (Rocky River) = 26.80 + 18.90 + 7.97 = $53.67

Question 7.

Gross margin per unit = Selling price per unit - Production cost per unit

Gross margin per unit (Sandy Beach) = 83.80 - 36.66 = $47.14

Gross margin per unit (Rocky River) = 105.00 - 53.67 = $51.33

Question 8.

Gross margin of Sandy Beach Model under traditional costing is $45.86 while that of under ABC costing method $47.14.

Gross margin of Rocky River Model under traditional costing is $53.02 and that of under ABC costing is $51.33